OpenStudy (anonymous):

factor the trinomials using the ac-method. s^2 - 6s + 3

OpenStudy (anonymous):

for using ac-method we need to first know what are the values of a, b and c a is the coefficient of the term with variable having power 2 b is the coefficient of the term with variable having power 1 c is the constant In your question s^2 - 6s + 3 a = 1 b = -6 c = 3 so ac = 1*3 = 3 we have to find factors of 3 whose product is 3 but sum is -6 the only factors of 3 are 1 and 3 and they cannot add u to -6 so this question cannot be factorised using ac method

OpenStudy (anonymous):

*cannot add up to -6